# How transparent is graphene? A surface science perspective on remote epitaxy

### v1

@@ V Outlook
How transparent is graphene? Our critical analysis of DFT slab calculations [5, 24] and analytical modeling [34] suggests that for weakly interacting graphene/substrate systems, the remote potential of the substrate is significantly attenuated by graphene. For many commonly used substrates, the remote substrate potential and the graphene potential have similar magnitude. We proposed a Fourier analysis to analyze the interference patterns between graphene and substrate in a bias-free way.
The literature on graphene-induced reconstructions suggest that systems with strong graphene/substrate interactions are a significant deviation from the transparent limit. In these systems, deviations from ideal bonding and variations in local registry of the graphene lattice on the substrate lattice can significantly impact sticking and template long range ordering of adatoms. This epitaxial ordering stretches the concept of remote epitaxy: here the graphene is not a passive transparent layer. Instead, the substrate imprints a deformation or local symmetry breaking in the graphene, which then templates the ordering of adatoms on the graphene. This can be viewed as a series of direct (nearest neighbor) substrate-graphene and graphene-film interactions, rather than a purely “remote” interaction that is transmitted through passive graphene. These effects are well studied in the limit of a few monolayers of adatoms on highly interacting graphene, and further studies are required to connect these insights to continuous films on graphene.
@@
Direct measurements of the remote potential, to test model predictions and distinguish effects of the remote potential versus growth kinetics. AFM-based force-distance spectroscopy [27, 83, 84, 28] and thermal desorption spectroscopy [85, 86, 87] are two possible methods to measure covalent and van der Waals contributions. Photoemission electron microscopy (PEEM) and low energy electron micrsocopy (LEEM) [88, 89, 90], Kelvin probe microsoppy (KPFM) [91, 92], and ballistic electron emission microscopy (BEEM) [93, 94] are suited for measuring electrostatic contributions.
@@
Systematic growth studies that vary the kinetics. Growth temperature, growth rates, precursor type, and defect identity and density are all key parameters that may enable tuning of remote epitaxy versus other mechanisms [19].
@@
To provide relevant information, we suggest that future first-principles theoretical calculations of remote epitaxy address these issues as well. Specifically, they must provide isolated slab calculations which include graphene layer(s) above the substrate, any relevant reconstructions of the graphene/substrate, and interaction between the substrate and film. Ideally, these calculations would capture information relevant to the kinetics of growth such as diffusion barriers for individual atoms or sliding barriers for islands. While requiring significant additional computation, calculations which also include graphene pinholes, defects, and step edges would be invaluable for understanding how these impact real-world bonding potentials and growth.
VI Acknowledgments

### v2

@@ VI Outlook
How transparent is graphene? Our critical analysis of DFT slab calculations [18, 33] and analytical modeling [44] suggests that for weakly interacting graphene/substrate systems, the remote potential of the substrate is significantly attenuated by graphene. For many commonly used substrates, the remote substrate potential and the graphene potential have similar magnitude. We proposed a Fourier analysis to analyze the interference patterns between graphene and substrate in a bias-free way. The literature on graphene-induced reconstructions suggest that systems with strong graphene/substrate interactions are a significant deviation from the transparent limit. In these systems, deviations from ideal bonding and variations in local registry of the graphene lattice on the substrate lattice can significantly impact sticking and template long range ordering of adatoms. These effects are well studied in the limit of a few monolayers of adatoms on highly interacting graphene, and further studies are required to connect these insights to continuous films on graphene.
@@
Direct measurements of the remote potential, to test model predictions and distinguish effects of the remote potential versus growth kinetics. AFM-based force-distance spectroscopy [36, 96, 97, 37] and thermal desorption spectroscopy [98, 99, 100] are two possible methods to measure covalent and van der Waals contributions. Photoemission electron microscopy (PEEM) and low energy electron microscopy (LEEM) [101, 102, 103], Kelvin probe microscopy (KPFM) [104, 105], and ballistic electron emission microscopy (BEEM) [106, 107] are suited for measuring electrostatic contributions.
@@
Explicit pairwise potential calculations that consider both the identity of the film and the substrate. Bonding is inherently a pairwise interaction, where changing the identity of one species in the pair changes that interaction. For example, Ga-Ga bonding is different than Ga-As bonding. However, with the exception of the kinetics literature [74, 75, 76, 77, 78, 79, 80, 81], most calculations of the remote potential to date[1, 18, 7] consider only the identity of the substrate, and do not explicitly consider the identity of the film. The film is only considered afterwards, e.g. via the adjustable scale factor in Eq 1 [33]. A true understanding of the remote potential requires explicit consideration of the identity of both film and substrate.
@@
Systematic growth studies that vary the kinetics. Growth temperature, growth rates, precursor type, and defect identity and density are all key parameters that may enable tuning of remote epitaxy versus other mechanisms [30]. The question of whether remote epitaxy “works” for a particular film / graphene (or other 2D material) / substrate system may not have a simple binary answer. Rather, it should also depend on what pathways are kinetically accessible.
4.
@@
To provide relevant information, we suggest that future first-principles theoretical calculations of remote epitaxy address these issues as well. Specifically, they must provide isolated slab calculations which include graphene layer(s) above the substrate, any relevant reconstructions of the graphene/substrate, and different substrate and film atoms, separated by enough vacuum that the top of the film does not interact with the bottom of the substrate. Ideally, these calculations would capture information relevant to the kinetics of growth such as diffusion barriers for individual atoms or sliding barriers for islands. While requiring significant additional computation, calculations which also include graphene pinholes, defects, and step edges would be invaluable for understanding how these impact real-world bonding potentials and growth.
VII Acknowledgments